# The Structure of Financial Equity Research Reports – Identification of the Most Frequently Asked Questions in Financial Analyst Reports to Automate Equity Research Using Llama 3 and GPT-4

### v1

@@ Abstract
The study analyzes 72 ERRs sentence-by-sentence, classifying their 4940 sentences into 169 unique question archetypes. We did not predefine the questions but derived them solely from the statements in the ERRs. This approach provides an unbiased view of the content of the observed ERRs. Subsequently, we used public corporate reports to classify the questions’ potential for automation. Answers were labeled “text-extractable” if the answers to the question were accessible in corporate reports.
78.7% of the questions in ERRs can be automated. Those automatable question consist of 48.2% text-extractable (suited to processing by large language models, LLMs) and 30.5% database-extractable questions. Only 21.3% of questions require human judgment to answer.
We empirically validate, using Llama-3-70B and GPT-4-turbo-2024-04-09 that recent advances in language generation and information extraction enable the automation of approximately 80% of the statements in ERRs. Surprisingly, the models complement each other’s strengths and weaknesses well.
@@
The full question list, including the archetypes and their frequency, will be made available online after peer review.
Keywords:
applied natural language processing text generation question answering generative AI natural language processing financial text
Working Paper
1 Introduction
@@
This empirical study aims to estimate the automation potential when writing equity research reports (ERRs) by systematically classifying statements of equity researchers into categories, and subsequently evaluating which sections and statements in ERRs require human judgment. One of the key contributions is a question list that analysts answer in ERRs. This question list gives a holistic overview of the ERR topics. We classify each question’s automation potential by comparing the data sources to the statements to see whether an automated system could have written the ERRs. If the share of automatable questions is high, this study may indicate that the automation of large parts of ERRs is feasible.
@@
ERRs are written mainly by sell-side banks and specialized research companies. The purpose of ERRs is ultimately to serve as a buy, hold, or sell recommendation (Asquith et al.,, 2005, p. 246). Two of the most common differences among ERRs are their content and intention. The content of ERRs can be to initiate a company’s coverage, provide an ordinary update to an already covered company, or provide an extraordinary update about a company. Subsequent recommendations to buy, hold, or sell may change or stay the same. One common update cycle for ERRs is quarterly ordinary updates. ERRs thus often contain recent facts from quarterly company-provided financial reports. Half of issued new ERRs generally fall into this category, being released closely after the company published new information (Asquith et al.,, 2005, p. 247). We review the literature on ERRs in section 2.1, discussing their importance, accuracy, and existing approaches for automation.
@@
We created the question list by manually reading each sentence of 72 ERRs. We mapped each phrase in the reports to a question. When we encountered a new question, we added it to the list. The result is a histogram of question occurrences (figure 2). We provide more detail about the methodology in section 4.
@@
This approach aligns with prior research (Asquith et al.,, 2005, p. 251), with the main difference being that we did not assume specific data fields (or questions) to be present in the reports ex-ante. Instead, we recorded each statement, derived a question from it, and counted the number of occurrences of each question. This approach ensures maximum unbiasedness in representing the landscape of ERRs.
@@
No systematic reviews of ERR automation exist. News feeds for financial news are already partially generated by AI systems. However, the automation of ERRs is not yet widely automated, albeit being feasible (Coleman et al.,, 2022). Some consumer-grade analyst houses such as Zacks.com use template-based automation to update their company profiles and overview articles. Longer texts, however, are still written by humans. ERRs are one example of such longer texts, and they are the focus of this study.
2 Literature
2.1 Review of the Financial Economics Literature
@@
Considering how strongly stock market prices react after ERRs are published, their importance for investors and, by extension, stock-listed corporations is evident in the literature (Bjerring et al.,, 1983; Elton et al.,, 1986; Liu et al.,, 1990; Beneish,, 1991; Stickel,, 1995).
@@
Womack, (1996) observed that analysts could predict the directionality of stock returns six months into the future (pp. 139, 163–165), indicating significant information content in ERRs. At least directionally, not necessarily with respect to the accuracy of the price targets, ERRs thus exhibit predictive accuracy higher than what can be expected from random guessing.
@@
A persistent problem with the reliability of ERRs is the reluctance of analysts to present negative recommendations, as Barber et al., (2001) and Mikhail et al., (2004) demonstrated. Asquith et al., (2005) report a rate of sell recommendations of only 0.5% in an ERR sample from 1997 to 1999 (p. 255), while Womack, (1996) report a rate of sell recommendations of 14% (p. 164) between 1989 and 1991. Prior research by Michaely and Womack, (1999) concludes that the reason for the low number of sell recommendations is likely that the companies covered by many financial analysts are the banks’ clients. Therefore, incentives arise to report too positively.
@@
According to a study by Asquith et al., (2005), more than half (54%) of analyst reports set price targets that are achieved within a year (pp. 278–279). This accuracy rate appears relatively low, considering that the majority of ERRs tend to offer conservative recommendations, with price targets slightly exceeding current prices, on average (Asquith et al.,, 2005, p. 256). The authors did not judge whether the analysts’ success rate was good or poor. In contrast to this outcome, Bradshaw et al., (2013) found that only 38% of analysts’ price targets are met within a one-year horizon (pp. 953–954). Bonini et al., (2010) provide another critical analysis of financial analysts’ accuracy (pp. 1177, 1193–1196, 1208). The unequivocal findings potentially strengthen the argument for automating ERRs to achieve more robust price targets (Gleason et al.,, 2013, pp. 80–81, for further discussion on report accuracy).
@@
Research by Coleman et al., (2022) shows that these issues can be mitigated by using a hybrid machine-human approach. The study presents a computer-aided approach that better balances the buy, hold, and sell recommendation frequencies, achieves better portfolio performance, and reduces the time required for writing reports.
@@
The previously presented studies show that the importance and accuracy of ERRs are studied extensively. Furthermore, there is extensive research on how stock performance is predicted by systematic factors, notably factor models by Eugene Fama and Kenneth French Fama and French, (1992, 1995, 2018, 2015) and by other studies from the field of risk factors and asset pricing (Carhart,, 1997; Asness et al.,, 2013).
@@
Asquith et al., (2005) performed a similar empirical analysis as our study. They ex-ante determined 30 variables of interest that were extracted from ERRs. There are differences in the objective, the data, and the methodology when comparing Asquith et al., to our study. They examined the accuracy of ERRs, how much they impact markets, and how independent research providers compare to sell-side banks. The data is from 1997 to 1999, and they selected only a subset of high-performing analysts. Also, they derived the 30 data fields from the objective before reading the ERRs; in this study, however, we read the ERRs and generated the questions ad-hoc when encountering new questions.
@@
2.2 Review of the Technical Literature
@@
Question answering (QA) systems commonly use text retrieval. Text retrievers get a question or a topic as input and search for fitting segments from a knowledge base. The retrieved span of text can be used directly as the answer or can be postprocessed by another system. Text retrieval from financial statements, even in combination with text from table captions, is already implemented by Chen et al., (2021). As retrievers are capable of choosing facts from knowledge bases of practically unlimited size, they are a key lower-level component needed to enable fact-based automated writing Borgeaud et al., (2021). It thus remains a core technology for automated ERR writing, at least until large language models’ context sizes become larger than they are today (Chen et al.,, 2023).
@@
As the stated goal of this study is to pave the way for automating the writing of ERRs, we provide a brief overview of the QA literature. QA is a subfield of natural language processing (NLP). It can be one of the domains that facilitate the automation of ERR writing.
@@
Dense passage retrieval uses latent representation of the question to search for an answer in a large corpus of text Karpukhin et al., (2020). Combining such a retrieval mechanism with a generative language model by including the retrieval outputs to the language model prompt, one gets a retrieval-augmented generator (RAG) as presented by Lewis et al., (2020). RAG continues to receive attention from the research community, as follow-up research on the topic shows (Izacard et al.,, 2022; Guu et al.,, 2020; Borgeaud et al.,, 2021). Its effectiveness in writing factually correct and fluent text makes RAG a technology that may facilitate the automation of ERR writing.
@@
With the emergence of the mentioned RAG methods, generative models are better capable of performing QA tasks. Furthermore, as models and their training data scale, their ability to store knowledge in their parameters increases, making them capable QA models even without external knowledge bases, as GPT-3 Brown et al., (2020) and the Llama models Touvron et al., (2023) demonstrate. In addition, advances in the expansion of the context length (Press et al.,, 2022; Chen et al.,, 2023) make it possible to provide more world knowledge into prompts.
@@
Figure 2: Histogram of Category Frequency, Grouped by Question Category. The histogram shows how frequently different information is displayed, grouped by question category. ERRs contain the financial category most frequently and usually display this category in text-or-tabular format or in text-tabular-or-graphical or tabular format. Company and market information are commonly in text-only format. Notably, stock information is rarely in text-only format. Only 2.82% of questions (weighted by their number of occurrences) were exclusively displayed in graphical format.
@@
As the general capabilities of language models grow, finance-specific language models also improve. Wu et al., (2023) develop a language model capable of understanding the nuances of a financial text. About half of their training data is finance-specific. The model performs well on financial QA (pp. 31–32). These generative language model developments increase the capabilities of state-of-the-art language models to generate factually correct ERRs.
3 Data
@@
We downloaded the ERRs dated from 2018 to 2023 from Bloomberg and Refinitiv Eikon. Each of the 72 reports had an average of seven pages, with the median being 6.8 pages per report. The shortest report is a one-pager, and the longest report has 20 pages. See also figure 8 for a histogram of page counts. We analyzed 493 pages across all reports. These statistics are in line with findings of previous research by (Asquith et al.,, 2005, p. 252). We sourced the ERRs from 23 different research providers, each contributing between 1 and 16 reports (figure 3).
@@
The number of statements per report ranges from nine to 115. The average is 68, and the median is 69. See also figure 1 for a histogram of statement counts. In sum, we analyzed 4940 statements (sentences).
4 Methods
@@
4.1 Question List and Question Categories
@@
We grouped the resulting 169 questions into five categories. The categories are: Financials, Company, Product, Stock, Market, Analysis. The questions are listed in tables 4-7.
4.2 Question Classification
@@
    Extractability: If the answer to a question is found in a corporate report, the question is extractable. To validate that our classification of extractability is correct, we ran the open-source Llama-3-70B model and the closed-source GPT-4-turbo-2024-04-09 model on 200 example questions, showing that these language models can indeed extract the answers to the questions when provided with annual reports as the prompt context. We describe this validation step in 4.4.
@@
The aggregated results are in table 2 and the original flags are in tables 4-7. We used company-issued reports to check whether the questions are extractable.
The annotation process required two iterations of reading through the reports. In the first reading process, the question list was created. In the second iteration, we labeled the extractability and the modality columns. The second iteration involved reading the ERRs again, and finding the source for the answer to each question. If a direct answer was matched in a single source, the extractability was marked as “extractable.” If the answer was not found in a single passage, the extractability was marked as “non-extractable.”
@@
4.3 Qualitative Validation of the Question List Using Expert Interviews and Prior Research
We conducted interviews with ten financial analysts. Each interview lasted 45 minutes. The interviews aimed to validate the relevance of our results and confirm whether the count-based ranking of questions was realistic. In the interviews, we were reassured that our results align with what would be expected in practice when considering the most important questions in ERRs. In some interviews, however, it became clear that financial analysts consider the management qualifications and subjective impressions about the competence of managers. This aspect is not captured in this study and will be hard to capture for automated systems.
In addition to validation through qualitative interviews, prior research by Asquith et al., (2005, p. 246) confirms that the top questions identified in our study match their findings about the most frequent ERR contents.
4.4 Validation of the Automation Potential by Comparing Human, Llama-3-70B, and GPT-4-turbo-2024-04-09 Report Generation Performance
We validate the claims made in this paper about the automation potential of specific questions by automating parts of the report generation. We test which questions are the hardest questions for language models to answer, which informs our assessment of the question “extractability.“
We use the open-source Llama-3-70B model by Meta AI and the closed-source GPT-4-turbo-2024-04-09 model by OpenAI. We set the temperatures of both models to zero (giving the models the chance to always use their true best guesses) and do not limit the number of output tokens. In cases where the context lengths of the models were exceeded, we split the contexts and concatenated the models’ outputs.
Figure 5: Share of Correct, Incorrect, Database-Extractable, and Non-Extractable Questions for GPT-4-turbo-2024-04-09 and Llama-3-70B. Stock-related questions can only be answered using financial market data. Those thus fall under a separate category that is automatable, yet not by using language models. Llama 3 has a slight edge over GPT-4 as it answered more questions correctly than GPT-4 did.
Our results show that Llama-3-70B is able to extract information from annual reports in 27% of the 200 questions. GPT-4 is able to extract the correct answer in 26% of the questions. The results are in line with our expectations and confirm that the models can indeed extract the answers to the questions when provided with annual reports as the prompt context.

### v2

# Identification of the Most Frequently Asked Questions in Financial Analyst Reports to Automate Equity Research Using Llama 3 and GPT-4

@@ Abstract
The study analyzes 72 ERRs sentence-by-sentence, classifying their 4964 sentences into 169 unique question archetypes. We did not predefine the questions but derived them solely from the statements in the ERRs. This approach provides an unbiased view of the content of the observed ERRs. Subsequently, we used public corporate reports to classify the questions’ potential for automation. Answers were labeled “text-extractable” if the answers to the question were accessible in corporate reports.
75.15% of the questions in ERRs can be automated using text extraction from text sources. Those automatable questions consist of 51.91% text-extractable (suited to processing by large language models, LLMs) and 24.24% database-extractable questions. Only 24.85% of questions require human judgment to answer.
We empirically validate, using Llama-3-70B and GPT-4-turbo-2024-04-09, that recent advances in language generation and information extraction enable the automation of approximately 80% of the statements in ERRs. Surprisingly, the models complement each other’s strengths and weaknesses well, indicating strong ensemble potential.
@@
The full question list, including the archetypes and their frequency, are available online (janspoerer.github.io/pop-spoerer-2025-financial-report-data).
Index Terms:
natural language processing (NLP), financial text, equity research reports, information extraction
I Introduction
@@
This study evaluates the automation potential of equity research reports by classifying analyst statements into categories and identifying which report components require human judgment. One of the key contributions is a question list that analysts answer in ERRs. This question list gives a holistic overview of ERR topics. We classify each question’s automation potential by comparing the data sources to the statements to see whether an automated system could have written the ERRs. If the share of automatable questions is high, this study may indicate that the automation of large parts of ERRs is feasible.
@@
ERRs are written mainly by sell-side banks and specialized research companies. The purpose of ERRs is ultimately to serve as a buy, hold, or sell recommendation [1, p. 246]. Two of the most common differences among ERRs are their content and intention. The content of ERRs can be to initiate a company’s coverage, provide an ordinary update to an already covered company, or provide an extraordinary update about a company. Subsequent recommendations to buy, hold, or sell may change or stay the same. ERRs often follow a specific update frequency. One common update cycle for ERRs is quarterly ordinary updates. ERRs thus often contain recent facts from quarterly company-provided financial reports. Half of all newly issued ERRs fall into this category, being released closely after the company published new information [1, p. 247].
We review the literature on ERRs in section II-A, discussing their importance, accuracy, and existing approaches for automation.
@@
We created the question list by manually reading each sentence of 72 ERRs. We mapped each phrase in the reports to a question. When we encountered a new question, we added it to the list. The result is a histogram of question occurrences (figure 2). We provide more detail about the methodology in section IV.
@@
This approach aligns with prior research [1, p. 251], with the main difference being that we did not assume specific data fields (or questions) to be present in the reports ex-ante. Instead, we recorded each statement, derived a question from it, and counted the number of occurrences of each question. This approach ensures maximum unbiasedness in representing the landscape of ERRs.
@@
No systematic reviews of ERR automation exist. News feeds for financial news are already partially generated by AI systems. However, ERR writing is not yet widely automated, albeit being feasible [2]. Some consumer-grade analyst houses such as Zacks.com use template-based automation to update their company profiles and overview articles. Longer texts, however, are still written by humans. ERRs are one example of such longer texts, and they are the focus of this study.
II Literature
II-A Review of the Financial Economics Literature
@@
Considering how strongly stock market prices react after ERRs are published, their importance for investors and, by extension, stock-listed corporations is evident in the literature [3, 4, 5, 6, 7].
@@
[8] observed that analysts could predict the directionality of stock returns six months into the future (pp. 139, 163–165), indicating significant information content in ERRs. At least directionally, not necessarily with respect to the accuracy of the price targets, ERRs thus exhibit predictive accuracy higher than what can be expected from random guessing.
@@
A persistent problem with the reliability of ERRs is the reluctance of analysts to present negative recommendations, as [9] and [10] demonstrated. [1] report a rate of sell recommendations of only 0.5% in an ERR sample from 1997 to 1999 (p. 255), while [8] report a rate of sell recommendations of 14% (p. 164) between 1989 and 1991. Prior research by [11] concludes that the reason for the low number of sell recommendations is likely that the companies covered by many financial analysts are the banks’ clients. Therefore, incentives arise to report too positively.
@@
According to a study by [1], more than half (54%) of analyst reports set price targets that are achieved within a year (pp. 278–279). This accuracy rate appears relatively low, considering that the majority of ERRs tend to offer conservative recommendations, with price targets slightly exceeding current prices, on average [1, p. 256]. The authors did not judge whether the analysts’ success rate was good or poor. In contrast to this outcome, [12] found that only 38% of analysts’ price targets are met within a one-year horizon (pp. 953–954). [13] provide another critical analysis of financial analysts’ accuracy (pp. 1177, 1193–1196, 1208). The unequivocal findings potentially strengthen the argument for automating ERRs to achieve more robust price targets [14, pp. 80–81, for further discussion on report accuracy].
@@
Research by [2] shows that these issues can be mitigated by using a hybrid machine-human approach. The study presents a computer-aided approach that better balances the buy, hold, and sell recommendation frequencies, achieves better portfolio performance, and reduces the time required for writing reports.
@@
The previously presented studies show that the importance and accuracy of ERRs are studied extensively. Furthermore, there is extensive research on how stock performance is predicted by systematic factors, notably factor models by [15, 16, 17, 18] and by other studies from the field of risk factors and asset pricing [19, 20].
@@
[1] performed a similar empirical analysis as our study. They ex-ante determined 30 variables of interest that were extracted from ERRs. There are differences in the objective, the data, and the methodology when comparing [1] to our study. They examined the accuracy of ERRs, how much they impact markets, and how independent research providers compare to sell-side banks. The data is from 1997 to 1999, and they selected only a subset of high-performing analysts. Also, they derived the 30 data fields from the objective before reading the ERRs; in this study, however, we read the ERRs and generated the questions ad-hoc when encountering new questions.
@@
[21] recently showed that the large language model GPT-4 can slightly outperform a simpler neural network in processing equity risk factors for stock analysis. They also show that humans are significantly worse at predicting stock returns than the authors’ language model-based system. This finding is in line with the aforementioned biases of human analysts.
II-B Review of the Technical Literature
@@
Text retrievers get a question or a topic as input and search for fitting segments from a knowledge base. The retrieved span of text can be used directly as the answer or can be postprocessed by another system. Text retrieval from financial statements, even in combination with text from table captions, is already implemented by [22], and earlier research on business text extraction has been a field of interest for years [23]. As retrievers are capable of choosing facts from knowledge bases of practically unlimited size, they are a key lower-level component needed to enable fact-based automated writing [24]. It thus remains a core technology for automated ERR writing, at least until large language models’ context sizes become larger than they are today [25].
@@
As the stated goal of this study is to pave the way for automating the writing of ERRs, we provide a brief overview of the question answering (QA) literature. QA is a subfield of NLP. It can be one of the domains that facilitate the automation of ERR writing.
@@
Dense passage retrieval uses a latent representation of a question to search for an answer in a large corpus of text [26]. Combining such a retrieval mechanism with a generative language model by including the retrieval outputs to the language model prompt, one gets a retrieval-augmented generator (RAG) as presented by [27]. RAG continues to receive attention from the research community, as follow-up research on the topic shows [28, 24, 29]. Its effectiveness in writing factually correct and fluent text makes RAG a technology that may facilitate the automation of ERR writing.
@@
With the emergence of the mentioned RAG methods, generative models are better capable of performing QA tasks. Furthermore, as models and their training data scale, their ability to store knowledge in their parameters increases, making them capable QA models even without external knowledge bases, as GPT-3 [30] and the Llama models [31] demonstrate. In addition, advances in the expansion of the context length [32, 25] make it possible to provide more world knowledge into prompts.
@@
Figure 2: Histogram of Category Frequency, Grouped by Question Category. The histogram shows how frequently different information is displayed, grouped by question category. ERRs contain the Financials category most frequently and usually display this category in text-or-tabular format or in text-tabular-or-graphical or tabular format. Company and Market information are commonly in text-only format. Notably, Stock information is rarely in text-only format. Only 2.82% of questions (weighted by their number of occurrences) were exclusively displayed in graphical format.
@@
As the general capabilities of language models grow, finance-specific language models also improve. [33] develop a language model capable of understanding the nuances of a financial text. About half of their training data is finance-specific, and model performs well on financial QA (pp. 31–32). These generative language model developments increase the capabilities of state-of-the-art language models to generate factually correct ERRs.
III Data
@@
We downloaded 72 ERRs dated from 2018 to 2023 from Bloomberg and Refinitiv Eikon. Each report had an average of seven pages, and the median is 6.8 pages per report. The shortest report is a one-pager, and the longest report has 20 pages. We analyzed 493 pages across all reports. These statistics are in line with findings of previous research by [1, p. 252]. We sourced the ERRs from 23 different research providers, each contributing between 1 and 16 reports.
@@
The number of statements per report ranges from nine to 115. The average is 68, and the median is 69. See also figure 1 for a histogram of statement counts. In sum, we analyzed 4964 statements (sentences).
IV Methods
@@
IV-A Question List and Question Categories
@@
We grouped the resulting 169 questions into five categories. The categories are: Financials, Company, Product, Stock, Market, Analysis.
IV-B Question Classification
@@
    Extractability: If the answer to a question is found in a corporate report, the question is extractable. To validate that our classification of extractability is correct, we ran the open-source Llama-3-70B and the closed-source GPT-4-turbo-2024-04-09 models on 200 example questions, showing that these language models can indeed extract the answers to the questions when provided with annual reports as the prompt context. We describe this validation step in section IV-D.
@@
The aggregated results are in table II. We used company-issued reports to check whether the questions are extractable.
The annotation process required two iterations of reading through the reports. In the first reading process, the question list was created. In the second iteration, we labeled the extractability and the modality columns. The second iteration involved reading the ERRs again, and finding the source for the answer to each question. If a direct answer was matched in a single text source, the extractability was marked as “extractable.” If the answer was found in a financial markets database (such as Bloomberg), the question was marked as “database-extractable.” If not found in single text passage or database, the extractability was marked as “non-extractable.”
@@
IV-C Qualitative Validation of the Question List Using Expert Interviews and Prior Research
As a sanity check, we conducted designated validation 45-minute interviews with ten financial analysts. In the interviews, we were reassured that our results align with what would be expected in practice when considering the most important questions in ERRs. In some interviews, however, it became clear that financial analysts consider the management qualifications and subjective impressions about the competence of managers. This aspect is not captured in this study and will be hard to capture for automated systems.
In addition to validation through qualitative interviews, prior research by [1, p. 246] confirms that the top questions identified in our study match their findings about the most frequent ERR contents.
IV-D Validation of the Automation Potential by Comparing Human, Llama-3-70B, and GPT-4-turbo-2024-04-09 Report Generation Performance
We validate the claims made in this paper about the automation potential of specific questions by automating those parts of the report generation that we have classified as “text-extractable.” We test which questions are the hardest questions for language models to answer, which informs our assessment of the question “text-extractability.”
We use the open-source Llama-3-70B model by [34] and the closed-source GPT-4-turbo-2024-04-09 model by [35]. We set the temperatures of both models to zero (giving the models the chance to always use their true best guesses) and do not limit the number of output tokens. In cases where the context lengths of the models were exceeded, we split the contexts and concatenated the outputs.
Figure 5: Share of Correct, Incorrect, Database-Extractable, and Non-Extractable Questions for GPT-4-turbo-2024-04-09 and Llama-3-70B, Weighted by Occurence. Stock-related questions can only be answered using financial market data. Those thus fall under a separate category that is automatable, yet not by using language models. Llama 3 has a slight edge over GPT-4 as it answered more questions correctly than GPT-4 did.
Our results show that Llama-3-70B is able to extract information from annual reports for 27% of the 200 questions. GPT-4 is able to extract the correct answer in 26% of the questions (figure 5). The results are in line with our expectations and confirm that the models can indeed extract the answers to the questions when provided with annual reports as the prompt context.